# Regret Analysis for Randomized Gaussian Process Upper Confidence Bound

### v1

@@ Abstract
Gaussian process upper confidence bound (GP-UCB) is a theoretically established algorithm for Bayesian optimization (BO), where we assume the objective function $f$ follows GP. One notable drawback of GP-UCB is that the theoretical confidence parameter $\beta$ increased along with the iterations is too large. To alleviate this drawback, this paper analyzes the randomized variant of GP-UCB called improved randomized GP-UCB (IRGP-UCB), which uses the confidence parameter generated from the shifted exponential distribution. We analyze the expected regret and conditional expected regret, where the expectation and the probability are taken respectively with $f$ and noises and with the randomness of the BO algorithm. In both regret analyses, IRGP-UCB achieves a sub-linear regret upper bound without increasing the confidence parameter if the input domain is finite. Finally, we show numerical experiments using synthetic and benchmark functions and real-world emulators.
@@ 1 Introduction
Bayesian optimization (BO) has become a popular black-box optimization method. BO aims to optimize with fewer function evaluations for the objective function, which is expensive to evaluate. For this purpose, BO sequentially evaluates the function value at the recommended input, which is chosen by maximizing an acquisition function based on a surrogate Bayesian model. This paper focuses on the BO using the Gaussian process (GP) surrogate model. GP-based BO has been widely used in many applications, such as automatic machine learning (Snoek et al., 2012), materials informatics (Ueno et al., 2016), and drug design (Korovina et al., 2020; Griffiths and Hernández-Lobato, 2020).
@@
GP upper confidence bound (GP-UCB) (Srinivas et al., 2010) is a seminal work for the regret guaranteed BO. Srinivas et al. (2010) showed the sub-linear regret bounds of GP-UCB, which become a fundamental technique for regret analysis in this field. On the other hand, GP-UCB suffers from the tuning of the confidence parameter, which is often written as $\beta_{t}$. The theoretical choice of the confidence parameter is too large and required to be increased. Hence, although the analysis of GP-UCB is solid and general, there is still an essential gap between theory and practice.
@@
To resolve this problem, Berk et al. (2020) have proposed the randomized variants of GP-UCB called RGP-UCB, where the confidence parameter follows the Gamma distribution. Although Berk et al. (2020) have shown the superior performance of randomized GP-UCB, their regret analysis appears to contain several technical issues (Appendix C of Takeno et al., 2023). We rectified the analysis and showed that GP-UCB with the randomized confidence parameter achieves the sub-linear expected regret under several conditions in the precedent conference version of this paper (Section 4.1 and Appendix D of Takeno et al., 2023). However, even in this analysis, the confidence parameter is still too large, as with GP-UCB.
@@
In this paper, we analyze the randomized GP-UCB using the confidence parameter generated from the shifted exponential distribution, referred to as improved randomized GP-UCB (IRGP-UCB). First, we analyze the expected regret, which evaluates the performance averaged by all the randomness of GP, the noises, and the BO algorithm itself. Our analysis tightens the expected regret upper bound, by which increasing the confidence parameter can be avoided if the input domain is finite. On the other hand, since the randomness of the BO algorithm is averaged, the expected regret analysis is not fair in comparing the expected regret upper bounds of the deterministic BO algorithm. That is, depending on the realization of the BO algorithm, the regret averaged by the randomness of GP and the noises may get worse. Therefore, we further analyze the conditional expected regret, where the expectation and the probability are taken respectively with GP and noises and with the randomness of the BO algorithm. Even in this analysis, IRGP-UCB does not require increasing the confidence parameter and maintains the same order of upper bounds.
In addition, we show that randomized GP-UCB can achieve the regret upper bounds with high probability by appropriately changing the distribution of the confidence parameters. However, in our analysis, the expectation of the confidence parameter must be increased to the same as GP-UCB. We will discuss the difficulty of showing high probability regret bounds of the randomized GP-UCB, which we leave as future work.
@@
This paper is organized as follows. Section 1.1 reviews the existing regret-guaranteed BO methods in the Bayesian setting, in which the objective function $f$ follows the GP. Section 2 provides the technical backgrounds. Section 3 shows the algorithm of the proposed method. In Section 4, we show the regret analyses for the expected and conditional expected regret. In addition, we discuss the high probability regret bounds of the randomized GP-UCB. Section 5 shows the wide range of numerical and real-world dataset experiments. Finally, Section 6 concludes.
@@ 1.1 Related Work
This study considers BO with the Bayesian setting, where the objective function $f$ is assumed to be a sample path from the GP. Various BO methods have been developed in the literature, for example, expected improvement (EI) (Mockus et al., 1978), entropy search (ES) (Hennig and Schuler, 2012), and predictive entropy search (PES) (Hernández-Lobato et al., 2014). Although the regret analysis of EI for the noiseless and frequentist setting, in which $f$ is an element of reproducing kernel Hilbert space, is provided in (Bull, 2011), the Bayesian setting has not been considered. Further, although the practical performance of ES and PES has been shown repeatedly, their regret analysis is an open problem.
@@
GP-UCB is one of the prominent studies for theoretically guaranteed BO methods. Srinivas et al. (2010) showed the high probability bound for cumulative regret. Furthermore, GP-UCB achieves the sub-linear expected regret upper bounds (Theorem 2.1 and B.1 of Takeno et al., 2023). However, the confidence parameter must be scheduled as $\beta_{t}\propto\log t$ in both regret analyses. Although many studies (for example, Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Janz et al., 2020) considered the frequentist setting, this study concentrates on the Bayesian setting. On the other hand, truncated variance reduction (TRUVAR) (Bogunovic et al., 2016) achieves high probability regret bounds. TRUVAR is based on the variance reduction in potential maximizers determined by UCB and lower confidence bounds. Hence, TRUVAR may be improved using randomization similar to GP-UCB.
@@
Another well-known regret-guaranteed BO method is Thompson sampling (TS) (Russo and Van Roy, 2014; Kandasamy et al., 2018; Takeno et al., 2024). TS achieves the sub-linear BCR bounds by sequentially observing the optimal point of the GP’s posterior sample path. Although TS does not require any hyperparameter, TS is often deteriorated by over-exploration, as discussed in (Shahriari et al., 2016). Recently, another posterior sampling-based BO called probability improvement from the maximum of sample path (PIMS) has been proposed. Known-best expected regret analyses of TS and PIMS are based on the proof technique of our precedent conference version of this paper (Takeno et al., 2023).
@@
Wang et al. (2016) have shown the regret analysis of the GP estimation (GP-EST) algorithm, which can be interpreted as GP-UCB with the confidence parameter defined using $\hat{m}$, an estimator of $\mathbb{E}[\max f(\boldsymbol{x})]$. Their analysis requires an assumption $\hat{m}\geq\mathbb{E}[\max f(\boldsymbol{x})]$, whose sufficient condition is provided in Corollary 3.5 of Wang et al. (2016). However, this sufficient condition does not typically hold, as discussed immediately after the corollary in Wang et al. (2016). Furthermore, the final convergence rate contains $\hat{m}$ itself, whose convergence rate is not clarified. In contrast, our Lemma 4.2 shows the bound for $\mathbb{E}[\max f(\boldsymbol{x})]$ under common regularity conditions. Wang and Jegelka (2017) have shown the regret analysis of max-value entropy search (MES). However, it is pointed out that their proof contains several technical issues (Takeno et al., 2022b).
@@
Another relevant literature is level set estimation (LSE) (Gotovos et al., 2013). LSE considers the classification problem of the expensive to evaluate black-box functions. After our preceding conference version of this paper (Takeno et al., 2023), a randomized version of LSE has been proposed by Inatsu et al. (2024b). Even for LSE, the randomization avoids increasing the confidence parameter in UCB-based LSE methods.
@@ 2 Background
In this section, we describe some background knowledge.
@@ 2.1 Bayesian Optimization
where $f$ is an black-box function, ${\cal X}\subset\mathbb{R}^{d}$ is an input domain, and $d$ is an input dimension. Here, black-box means that we can access only function values contaminated by noises. In addition, we consider the case that $f$ is expensive to evaluate. Therefore, BO sequentially observes the noisy function value with some policy so that the number of function evaluations is minimized as far as possible. This policy is determined as acquisition functions (AFs) $a:{\cal X}\rightarrow\mathbb{R}$, which quantify the utility of function evaluation. That is, at each iteration $t$, BO queries $\boldsymbol{x}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}a(\boldsymbol{x})$ and obtain $y_{t}=f(\boldsymbol{x}_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is an additive noise.
BO generally employs the GP surrogate model (Rasmussen and Williams, 2005) due to its flexible prediction capability and analytical computations. The GP surrogate model assumes that $f$ is a sample path from a GP with a zero mean and a kernel function $k:{\cal X}\times{\cal X}\mapsto\mathbb{R}$, and $\epsilon_{t}\sim{\cal N}(0,\sigma^{2})$ is i.i.d. Gaussian noise with a positive variance $\sigma^{2}>0$. At each iteration $t$, a dataset ${\cal D}_{t-1}\coloneqq\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{t-1}$ is already obtained from the nature of BO. Then, the posterior distribution $p(f\mid{\cal D}_{t-1})$ is a GP again. The posterior mean and variance of $f(\boldsymbol{x})$ are derived as follows:
@@ 2.2 Regularity Assumptions
We assume that the GP surrogate model is correct. That is, $\epsilon_{t}\sim{\cal N}(0,\sigma^{2})$ is i.i.d. Gaussian noise with a positive variance $\sigma^{2}>0$, and the underlying true function $f$ is a sample path from a known GP with a zero mean and a kernel function $k$. Here, zero mean is without loss of generality (Rasmussen and Williams, 2005). In addition, we assume that the kernel function $k$ is normalized. Therefore, the kernel $k(\boldsymbol{x},\boldsymbol{x}^{\prime})$ satisfies the condition $k(\boldsymbol{x},\boldsymbol{x}^{\prime})\leq 1$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in{\cal X}$.
@@
This assumption holds for stationary (Genton, 2001) and four times differentiable kernels (Theorem 5 of Ghosal and Roy, 2006), such as RBF kernel and Matérn kernels with a parameter $\nu>2$, as discussed in Section 4 of Srinivas et al. (2010). Therefore, this assumption is commonly used, for example, in Srinivas et al. (2010); Kandasamy et al. (2018); Paria et al. (2020); Takeno et al. (2024).
@@ 2.3 Regret
In this paper, we evaluate the performance of BO methods by cumulative regret (Srinivas et al., 2010; Russo and Van Roy, 2014; Kandasamy et al., 2018; Paria et al., 2020; Takeno et al., 2024). We incur the instantaneous regret $r_{t}=f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$ for each function evaluation. The cumulative regret for $T$ iterations is defined as the sum of instantaneous regrets over $T$ function evaluations defined as follows:
@@
Especially, the BO method, which achieves sub-linear cumulative regret, that is, $\lim_{T\rightarrow\infty}R_{T}/T=0$, is called no-regret. No-regret property also indicates that the best regret $\min_{t\in[T]}r_{t}$ converges to 00 since $\min_{t\in[T]}r_{t}<R_{T}/T$. Therefore, our main goal of the theoretical analysis is to show the sub-linearity of cumulative regret. Note that regret is a random quantity due to $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and the randomness of the BO algorithm.
@@
where $I$ is the Shanon mutual information.

### v3

@@ Abstract
Gaussian process upper confidence bound (GP-UCB) is a theoretically established algorithm for Bayesian optimization (BO), where we assume the objective function $f$ follows a GP. One notable drawback of GP-UCB is that the theoretical confidence parameter $\beta$ increases along with the iterations and is too large. To alleviate this drawback, this paper analyzes the randomized variant of GP-UCB called improved randomized GP-UCB (IRGP-UCB), which uses the confidence parameter generated from the shifted exponential distribution. We analyze the expected regret and conditional expected regret, where the expectation and the probability are taken respectively with $f$ and noise and with the randomness of the BO algorithm. In both regret analyses, IRGP-UCB achieves a sub-linear regret upper bound without increasing the confidence parameter if the input domain is finite. Furthermore, we show that randomization plays a key role in avoiding an increase in confidence parameter by showing that GP-UCB using a constant confidence parameter can incur linearly growing expected cumulative regret. Finally, we show numerical experiments using synthetic and benchmark functions and real-world emulators.
@@ 1 Introduction
Bayesian optimization (BO) has become a popular black-box optimization method. BO aims to optimize with fewer function evaluations for an objective function, which is expensive to evaluate. For this purpose, BO sequentially evaluates the function value at the recommended input, which is chosen by maximizing an acquisition function based on a surrogate Bayesian model. This paper focuses on BO using a Gaussian process (GP) surrogate model. GP-based BO has been widely used in many applications, such as automatic machine learning (Snoek et al., 2012), materials informatics (Ueno et al., 2016), and drug design (Korovina et al., 2020; Griffiths and Hernández-Lobato, 2020).
@@
GP upper confidence bound (GP-UCB) (Kushner, 1964; Srinivas et al., 2010) is a seminal work for regret-guaranteed BO. Srinivas et al. (2010) show the sub-linear regret bounds of GP-UCB, which has become a fundamental technique for regret analysis in this field. On the other hand, GP-UCB suffers from the tuning of the confidence parameter, which is often written as $\beta_{t}$. The confidence parameter for the theoretical analysis is too large and needs to increase over time. Hence, although the analysis of GP-UCB is solid and general, there is still an essential gap between theory and practice.
@@
To resolve this problem, Berk et al. (2020) propose the randomized variants of GP-UCB called RGP-UCB, where the confidence parameter follows the Gamma distribution. Although Berk et al. (2020) show the superior performance of randomized GP-UCB, their regret analysis appears to contain several technical issues (Appendix C of Takeno et al., 2023). We rectified the analysis and showed that GP-UCB with randomized confidence parameters achieves sub-linear expected regret upper bounds under several conditions in the previous conference version of this paper (Section 4.1 and Appendix D of Takeno et al., 2023). However, even in this analysis, the confidence parameter is too large and needs to increase over time, as with GP-UCB.
@@
In this paper, we analyze randomized GP-UCB using confidence parameters generated from the shifted exponential distribution, referred to as improved randomized GP-UCB (IRGP-UCB). First, we show sub-linear expected regret upper bounds of IRGP-UCB, where the regret is averaged over all the randomness of the GP, the noise, and the BO algorithm itself. In particular, our analysis reveals that IRGP-UCB can achieve a sub-linear expected regret without increasing the confidence parameter when the input domain is finite. On the other hand, since the randomness of the BO algorithm is averaged, the expected regret analysis is not fair in comparing the expected regret upper bounds of deterministic BO algorithms. That is, depending on the realization of the BO algorithm, the regret averaged by the randomness of GP and the noise may get worse. Therefore, we further analyze the conditional expected regret averaged over the randomness of the GP and the noise and conditioned on the randomness of the BO algorithm. Even in this analysis, IRGP-UCB does not require increasing the confidence parameter and maintains the same order of upper bounds. In addition, although the expectation of the randomized confidence parameters needs to increase, we show that randomized GP-UCB can achieve high-probability regret upper bounds, which implies that randomizing confidence parameters does not diminish the theoretical performance compared with the existing result (Srinivas et al., 2010).
The above expected and conditional expected regret analyses raise the question of whether randomization is necessary for achieving a sub-linear expected regret bound. We show an affirmative answer for this question by showing a problem instance where GP-UCB using a constant confidence parameter incurs linear expected cumulative regret even in finite input domains. Hence, at least for the expected regret analysis, we can see that randomizing the confidence parameters plays a key role in avoiding increasing confidence parameters.
@@
This paper is organized as follows. Section 1.1 reviews the existing regret-guaranteed BO methods in the Bayesian setting, in which the objective function $f$ follows the GP. Section 2 provides the technical background. Section 3 shows the algorithm of the proposed method. Section 4 shows the analyses for the expected regret upper bound of IRGP-UCB (Section 4.1), the conditional expected regret upper bound of IRGP-UCB (Section 4.2), the high-probability regret upper bound of randomized GP-UCB (Section 4.3), and the expected regret lower bound of GP-UCB with a constant confidence parameter on a specific problem instance (Section 4.4). Section 5 shows the wide range of numerical and real-world dataset experiments. Finally, Section 6 concludes.
Compared with the previous conference version of this paper (Takeno et al., 2023), we add the following analyses in addition to reorganizing the entire paper. First, we add the conditional expected regret analysis of IRGP-UCB in Section 4.2. Second, we show that randomized GP-UCB with increasing confidence parameters can achieve high-probability sub-linear cumulative regret bounds in Section 4.3. Third, in Section 4.4, we show that GP-UCB using a constant confidence parameter can incur linear expected cumulative regret. In addition, we have removed an unnecessary assumption, the kernel’s stationarity, from theorems and reorganized Lemma 4.2 (Lemma 4.2 of (Takeno et al., 2023)) to improve the lemma’s usefulness for the proof. Finally, we have added one baseline method called the probability of improvement from the maximum of sample path (Takeno et al., 2024) in experiments.
@@ 1.1 Related Work
The theoretical assumptions for BO are mainly twofold: the Bayesian setting (Srinivas et al., 2010; Russo and Van Roy, 2014; Kandasamy et al., 2018), where the objective function $f$ is assumed to be a sample path from a GP, and the frequentist setting (Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Janz et al., 2020), where the objective function $f$ is an element of a reproducing kernel Hilbert space (RKHS). Hence, we can construct credible intervals directly in the Bayesian setting since $f$ follows the posterior distribution. On the other hand, although $f$ does not follow the posterior distribution in the frequentist setting, confidence intervals for $f$ based on GP models have been developed (Srinivas et al., 2010; Chowdhury and Gopalan, 2017). Based on these tools, we can design BO algorithms that have regret guarantees using GP models. Note that since a sample path from a GP does not belong to the RKHS with bounded norm, neither analysis subsumes the other (See, e.g., the last paragraph in Section 4 of (Srinivas et al., 2010)). This study considers BO in the Bayesian setting.
@@
Various BO methods have been developed in the literature, for example, expected improvement (EI) (Mockus et al., 1978), entropy search (ES) (Hennig and Schuler, 2012), and predictive entropy search (PES) (Hernández-Lobato et al., 2014). The theoretical analysis of EI for the noiseless case, where function values are not contaminated by noise, has been performed in (Vazquez and Bect, 2010; Grünewälder et al., 2010; Bull, 2011) for both Bayesian and frequentist settings. However, the regret analysis of EI for the noisy case is still limited to the frequentist setting (Wang and de Freitas, 2014; Tran-The et al., 2022). Further, although the practical performance of ES and PES has been shown repeatedly, their regret analysis is an open problem.
@@
GP-UCB is one of the prominent studies for theoretically guaranteed BO methods. Srinivas et al. (2010) show the high-probability bound for cumulative regret. Furthermore, GP-UCB achieves the sub-linear expected regret upper bounds (Theorem 2.1 and B.1 of Takeno et al., 2023). However, the confidence parameter must be scheduled as $\beta_{t}\propto\log t$ in both regret analyses. Although many studies (for example, Srinivas et al., 2010; Chowdhury and Gopalan, 2017; Janz et al., 2020) considered the frequentist setting, this study concentrates on the Bayesian setting. On the other hand, truncated variance reduction (TRUVAR) (Bogunovic et al., 2016) achieves high-probability regret bounds. TRUVAR is based on the variance reduction in potential maximizers determined by UCB and lower confidence bounds. Hence, TRUVAR may be improved using randomization similar to GP-UCB.
@@
Another well-known regret-guaranteed BO method is Thompson sampling (TS) (Russo and Van Roy, 2014; Kandasamy et al., 2018; Takeno et al., 2024). TS achieves sub-linear BCR bounds by sequentially observing the optimal point of the GP’s posterior sample path. Although TS does not require any hyperparameter, TS often deteriorates by over-exploration, as discussed in (Shahriari et al., 2016). Recently, another posterior sampling-based BO called probability improvement from the maximum of sample path (PIMS) has been proposed. Known-best expected regret analyses of TS and PIMS are based on the proof technique of our previous conference version of this paper (Takeno et al., 2023).
@@
Wang et al. (2016) show regret analysis of the GP estimation (GP-EST) algorithm, which can be interpreted as GP-UCB with the confidence parameter defined using $\hat{m}$, an estimator of $\mathbb{E}[\max f(\boldsymbol{x})]$. Their analysis requires an assumption $\hat{m}\geq\mathbb{E}[\max f(\boldsymbol{x})]$, whose sufficient condition is provided in Corollary 3.5 of Wang et al. (2016). However, this sufficient condition does not typically hold, as discussed immediately after the corollary in Wang et al. (2016). Furthermore, the final convergence rate contains $\hat{m}$ itself, whose convergence rate is not clarified. In contrast, our Lemma 4.2 shows the bound for $\mathbb{E}[\max f(\boldsymbol{x})]$ under common regularity conditions. Wang and Jegelka (2017) show regret analysis of max-value entropy search (MES). However, it is pointed out that their proof contains several technical issues (Takeno et al., 2022b).
Scarlett (2018) shows that, under several conditions, there exists an algorithm that achieves a tighter high-probability regret upper bound compared with other well-known algorithms, such as GP-UCB and TS. However, the algorithm shown by Scarlett (2018) depends on the increasing confidence parameters for the construction of confidence bounds. Therefore, their algorithm also has the same gap between theoretical analysis and practical effectiveness as GP-UCB.
@@
Another relevant line of work is level set estimation (LSE) (Gotovos et al., 2013). LSE considers the classification problem of the expensive-to-evaluate black-box functions. After our preceding conference version of this paper (Takeno et al., 2023), a randomized version of LSE has been proposed by Inatsu et al. (2024b). Even for LSE, the randomization avoids increasing the confidence parameter in UCB-based LSE methods.
@@ 2.1 Bayesian Optimization
where $f$ is a black-box function, ${\cal X}\subset\mathbb{R}^{d}$ is an input domain, and $d$ is an input dimension. Here, black-box means that we can access only function values and cannot obtain other information, such as gradients. Furthermore, we assume that observations are contaminated by noise. In addition, we consider the case that $f$ is expensive to evaluate. Therefore, BO sequentially observes the noisy function value with some policy so that the number of function evaluations is minimized as far as possible. This policy is determined by an acquisition function (AF) $a:{\cal X}\rightarrow\mathbb{R}$, which quantifies the utility of function evaluation. That is, at each iteration $t$, BO queries $\boldsymbol{x}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}a(\boldsymbol{x})$ and obtains $y_{t}=f(\boldsymbol{x}_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is additive noise. Regularity assumptions of $f$ and $\epsilon_{t}$ are described in Section 2.2.
BO generally employs the GP surrogate model (Rasmussen and Williams, 2005) due to its flexible prediction capability and analytical computations. The GP surrogate model can be derived under assumptions that $f$ is a sample path from a GP with a zero mean and a kernel function $k:{\cal X}\times{\cal X}\mapsto\mathbb{R}$, and $\epsilon_{t}\sim{\cal N}(0,\sigma^{2})$ is i.i.d. Gaussian noise with a positive variance $\sigma^{2}>0$. At each iteration $t$, a dataset ${\cal D}_{t-1}\coloneqq\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{t-1}$ is already obtained from the nature of BO. Then, the posterior distribution $p(f\mid{\cal D}_{t-1})$ is a GP again. The posterior mean and variance of $f(\boldsymbol{x})$ are derived as follows:
@@ 2.2 Regularity Assumptions
We assume that the GP surrogate model is correct. That is, $\epsilon_{t}\sim{\cal N}(0,\sigma^{2})$ is i.i.d. Gaussian noise with a positive variance $\sigma^{2}>0$, and the underlying true function $f$ is a sample path from a known GP with a zero mean and a kernel function $k$. Here, the zero mean is without loss of generality (Rasmussen and Williams, 2005). In addition, we assume that the kernel function $k$ is normalized. Therefore, the kernel $k(\boldsymbol{x},\boldsymbol{x}^{\prime})$ satisfies the condition $k(\boldsymbol{x},\boldsymbol{x}^{\prime})\leq 1$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in{\cal X}$.
@@
This assumption holds for stationary (Genton, 2001) and four times differentiable kernels (Ghosal and Roy, 2006, Theorem 5 of), such as RBF kernel and Matérn kernels with a parameter $\nu][]2$, as discussed in Section 4 of Srinivas et al. (2010). Therefore, this assumption is commonly used, for example, in Srinivas et al. (2010); Kandasamy et al. (2018); Paria et al. (2020); Takeno et al. (2024).
@@ 2.3 Regret
In this paper, we evaluate the performance of BO methods by cumulative regret (Srinivas et al., 2010; Russo and Van Roy, 2014; Kandasamy et al., 2018; Paria et al., 2020; Takeno et al., 2024). We incur the instantaneous regret $r_{t}=f(\boldsymbol{x}^{*})-f(\boldsymbol{x}_{t})$ for each function evaluation. The cumulative regret for $T$ iterations is defined as the sum of instantaneous regrets over $T$ function evaluations, defined as follows:
@@
In particular, BO methods that achieve sub-linear cumulative regret (i.e., such that $\lim_{T\rightarrow\infty}R_{T}/T=0$) are called no-regret. No-regret property also indicates that the best regret $\min_{t\in[T]}r_{t}$ converges to 00 since $\min_{t\in[T]}r_{t}<R_{T}/T$. Therefore, the main goal of theoretical analysis is to show sub-linearity of cumulative regret. Note that regret is a random quantity due to $f$, $\{\epsilon_{t}\}_{t\geq 1}$, and the randomness of the BO algorithm.
@@
where $I$ is the Shannon mutual information.
@@
2.4 GP-UCB
GP-UCB selects the next input by maximizing UCB as follows:
$\boldsymbol{x}_{t}=\operatorname*{arg\,max}_{\boldsymbol{x}\in{\cal X}}\mu_{t-1}(\boldsymbol{x})+\beta^{1/2}_{t}\sigma_{t-1}(\boldsymbol{x}).$
In the existing regret analyses (Srinivas et al., 2010; Takeno et al., 2023), to obtain the sub-linear cumulative regret bounds, the confidence parameter $\beta_{t}$ must be scheduled as $\beta_{t}=\Theta\bigl{(}\log(|{\cal X}|t)\bigr{)}$ for finite input domains.